OpenStudy (anonymous):

Using each of the Integers (-2), (+8), (-6), and (+1) at least once, create four different expressions that include + - / and x at least once and result in an answer between (-10) and (+10) URGENT

OpenStudy (anonymous):

-2 * +1 / +6 + 8 - (-2) = 9 5/6

OpenStudy (anonymous):

thats one!

OpenStudy (anonymous):

thankyou!

OpenStudy (anonymous):

(-2)-(-6)/8*1+(-2) = - 1 1/2

OpenStudy (anonymous):

8 / 6 * 1 +(-2) -1 = -2 2/3